Sam lives on a lot that he thought was a square, 157 feet by 157 feet. When he had it surveyed, he discovered that one side was actually 2 feet longer than he thought and the other was actually 2 feet shorter than he thought. How much less area does he have than he thought he had?
Answer by checkley75(3666)

157*157=24,649 what he thought he had.
(157-2)(157+2)=155*159=24,645
24,649-24,645=4 square feet less than the original measurements.